OpenStudy (anonymous):

Write the explicit formula for the sequence with a1 = 120 and r = 0.3. Then generate the first five terms.

jimthompson5910 (jim_thompson5910):

the nth term is an = a1*r^(n-1) an = 120*0.3^(n-1)

jimthompson5910 (jim_thompson5910):

Each term is found by plugging in n = 1, n = 2, n = 3, etc

jimthompson5910 (jim_thompson5910):

Does that make sense?